How much does it cost to commute to work? How does distance of your commute relate to your fuel usage? Data were collected on distance between a persons home and primary work location (distance) and the persons monthly expenditures on gasoline for their vehicle (gas) for 17 individuals. The regression analysis is given below. One of the 17 individuals in the sample lived 12 miles from work and spent $149.25 on gas. What is this individuals residual value? Give your answer to 2 decimal places. For help on how to input a numeric answer, please see Instructions for inputting a numeric response Simple linear regression results: Dependent Variable: gas Independent Variable: distance gas 118.80432+2.1765497 distance Sample size: 17 R (correlation coefficient) 08107 R-sq 0.6571686 Estimate of error standard deviation: 17.030962 Parameter estimates: Parameter Estimate Std. Err. DF T-Stat P-Value Intercept 118.80432 9.172447 15 12.952303 0.0001 Slope 2.1765497 0.4059054 15 5.3622093 <0.0001 Answer:
Answer #1

From the given output, for x = 12:

y = 118.80432 + 2.1765497*12

y = 144.92

Hence,

Residual = 149.25 - 144.92 = 4.33